Differential Equations
Confused on how to do this
(2) (Spts) Find the explicit solution of the initial value problem 12 = 7e-2 – 2y where y(0) = 1.
Answer #1

consider the diff.eon dy =ner_zy le dj + 2y = xe * = dy + Pay = Q Cor) This is first order linear diff.eon. Hoa Pa)=3, Q1) =Its goneral solution is 9.(J.F) = (Q(x). If de to yees - Sae? 22 dhe te yezze - green de to gexe + C (San= uso-S (dusu) Lycx